OpenStudy (anonymous):

Solve the system. y = x + 2 and y = -x + 0 A. (1, -1) B. (-1, 1) C. (1, 3) D. (-1, 3) E. no solution

OpenStudy (anonymous):

B

jhonyy9 (jhonyy9):

y=x+2 y=-x so than x+2=-x so 2x=-2 so x= -2/2 =-1 than y=-(-1)=1 - so B is right